An ideal gas in a sealed container has an initial volume of 2.50 L. At constant pressure, it is cooled to19.00 ∘C, where its final volume is 1.75 L. What was the initial temperature?

?1=

∘C
Answer #1

1 AS Ni=2.5L T2 = lgic=273-15°ct 19 = 292.15k - V2 = 1.752 = ? As we know According to channes Law, 1 : Tiam Tz V2 Ti= 2.5L x